Construct the 99% confidence interval for the population mean duration of the longdistance calls from the town. 


1
Expert's answer
2021-11-03T15:05:38-0400

Assumptions:

mean "\bar{x}= 12"

sample standard deviation s= 2

Sample size n = 100

alpha = 1%

The critical value for α=0.01 and df= n-1= 99 is "t_c = 2.626"

"CI = \bar{x} ± t_c \times \frac{s}{\sqrt{n}} \\

CI = 12 ± 2.626 \times \frac{2}{\sqrt{100}} \\

CI = (11.475, 12.525)"